# Understanding Parents’ Perceptions and Practices Toward Children’s Security and Privacy in Virtual Reality

### v1

@@ Abstract
Recent years have seen a sharp increase in underage users of virtual reality (VR), where security and privacy (S&P) risks such as data surveillance and self-disclosure in social interaction have been increasingly prominent. Prior work shows children largely rely on parents to mitigate S&P risks in their technology use. Therefore, understanding parents’ S&P knowledge, perceptions, and practices is critical for identifying the gaps for parents, technology designers, and policymakers to enhance children’s S&P. While such empirical knowledge is substantial in other consumer technologies, it remains largely unknown in the context of VR. To address the gap, we conducted in-depth semi-structured interviews with 20 parents of children under the age of 18 who use VR at home. Our findings highlight parents generally lack S&P awareness due to the perception that VR is still in its infancy. To protect their children’s interaction with VR, parents currently primarily rely on active strategies such as verbal education about S&P. Passive strategies such as parental controls in VR are not commonly used among our interviewees, mainly due to their perceived technical constraints. Parents also highlight that a multi-stakeholder ecosystem must be established towards more S&P support for children in VR. Based on the findings, we propose actionable S&P recommendations for critical stakeholders, including parents, educators, VR companies, and governments.
@@ 1 Introduction
Furthermore, users including children have flooded social VR applications (e.g., Horizon Worlds, Rec Room, and VRChat) [5, 3, 4], posing more S&P risks. In internet-enabled social VR applications, besides data surveillance, users also face S&P risks in social interaction, including self-disclosure [13]. Prior work has suggested that social VR users easily disclose a wide range of privacy to strangers, including emotions, life experiences, voices, and even real-life appearances [13].
@@ 2 Related Work
Children’s S&P risks in VR. VR technologies collect a vast amount of sensitive data that has been categorized by Dick [7] into four types – (1) observable data that generate users’ virtual presence, e.g., avatars and real-time in-world interactions [28, 29]; (2) observed data that enhance immersive experiences, e.g., motion/eye-tracking [8, 9, 29, 30, 31]; (3) computed data that improve user services, e.g., user profiles for advertisements; and (4) associated data that allow Internet-enabled functions, e.g., IP addresses, usernames, and friend lists [32]. Unlike many other consumer technologies, VR collects these sensitive data that are indispensable for its core utility– immersive experiences [33]. Specifically, users have to trade their motion-tracking data for the full-body tracking experience; the data collected via external-facing cameras for detecting obstacles around; and voice data for social experiences [7]; just to name a few. Due to the unique immersive nature of VR technologies, researchers have been calling for policymakers to reform existing privacy laws and policies such as COPPA [34] and HIPAA [35] that would unnecessarily limit users’ experiences of VR technologies [7]. In addition, a 2018 study has revealed that VR developers considered VR privacy policies problematic for providing developers with few standards and guidelines regarding S&P [6].
@@
In the context of social interaction, more S&P risks have emerged as VR evolves to become a popular form of social technology [13, 36, 37]. Mainstream social VR platforms, including VRChat, Horizon Worlds, Rec Room, and more, allow users to socialize with others through head-mounted displays (HMDs) [38, 39]. In this vein, a volume of social VR work has identified numerous privacy and safety risks. Maloney et al. [13] have found that social VR users tend to feel comfortable and safe to share emotions and life experiences due to the anonymity afforded by social VR. In addition, users’ physical traits, such as their voices and real-life appearances, can also be disclosed since (1) voice chats are ubiquitous in social VR and (2) some users prefer to represent themselves with a virtual avatar close to their appearance in real life. More recently, with the increasing popularity of social VR streaming, users’ privacy has been violated by streamers who broadcast in public VR spaces [36].
@@
Additionally, due to the age-mixed environment [4, 3, 40], children who are considered part of the marginalized community in social VR often face more tailored risks, such as predators, harassment, cyberbullying, inappropriate adult content [5, 41, 42, 43, 44], etc. However, compared to such safety threats that have been widely studied, children’s S&P in social VR have been much more marginalized in scholarly work. Previous studies have collectively pointed out that bystanders and parents have concerns for children’s privacy as they interact with strangers in VR [4, 3]. Notably, one study has suggested VR privacy guidelines be made more straightforward to align with children’s reading and literacy levels [5]. To further provide actionable S&P-enhancing recommendations, more research centered around children’s S&P is needed. Since children heavily rely on parents’ decision-making of their online S&P [14, 45, 15, 46, 16], a critical gap exists in terms of parents’ attitudes and knowledge around children’s S&P in VR. Our work fills this gap by identifying and bridging parents’ knowledge gap in children’s S&P in VR.
@@
Parents’ risk mitigation strategies toward technologies at home. Nissenbaum defines privacy as “neither a right to secrecy nor a right to control, but a right to appropriate flow of personal information.” [47]. This definition is especially applicable to online social contexts where privacy risks arise from self-disclosure [48, 47, 49, 50, 51, 52]. When self-disclosure occurs online, privacy is about the appropriate flow of information [53]. With children’s significantly increased online activities [54, 55, 56], more unexpected information flows have raised new privacy concerns [16, 57]. For instance, information disclosed via email can be retained longer than face-to-face communications [16]. It is critical to protect children’s privacy online since (1) privacy is an important space for children to practice making decisions, create boundaries, and foster independence [48, 53]; (2) privacy can benefit children’s experience in identity play, relationship building, and peer communications [58, 16].
@@
Security is another important aspect of children’s interactions with technology. Prior work suggests three goals of information security [16, 59]: (1) maintaining personal information integrity when stored and transmitted through networks, (2) authenticating users with access to personal information, and (3) preserving personal information confidentiality. This work uses the above-mentioned definitions of S&P to unpack parents’ risk perceptions and mitigation strategies.
@@
Given the importance of S&P for children’s use of technology, parents have the responsibility of protecting children’s digital S&P at home [60, 61]. Kumar et al. [16] found that children aged 9 years old or younger were most likely to seek help from parents when confronted with unknown risks online and largely relied on parents’ explicit rules to deal with their digital S&P risks. Other research in this vein has suggested that children of different ages trust their parents’ S&P protections and guidance such as on data sharing [62, 63, 60]. Additionally, another line of research suggests that parents’ S&P attitudes and practices may shape that of children [60, 64, 21, 65, 17, 18, 19]. For example, Shin et al. [17] reported that parents who self-identified as more S&P-focused also tended to better direct children’s S&P behavior online.
@@
Parents adopt a variety of strategies to mitigate the risks of children’s technology use [16], including talking to children, viewing children’s accounts, using parental controls, being around their children when using technology, setting rules and boundaries for usage, etc. [66, 67, 20, 21]. Having examined different parental strategies, Wisniewski et al. [22] summarized three types of strategies in the Teen Online Safety Strategies (TOSS) framework: (1) passive monitoring, (2) rules and restrictions in usage, and (3) active mediation such as through having conversations with children. However, when mitigating the risks, parents face a number of challenges, such as underestimating the potential risks faced by children [68, 69, 23], undermining parent-child trust [23, 70, 71, 72], and an incomplete understanding of technologies compared to their children [23, 73, 24]. Therefore, it is crucial to investigate parents’ S&P practices around children’s technology use at home and identify the tradeoffs in their decision-making to help parents and technology designers work together towards safer and more inclusive technologies for children. Unfortunately, while such knowledge is substantial in other consumer technologies such as smartphones [25, 26] and Internet-enabled applications [16, 27], it is largely unknown in the VR space. Therefore, we aim to bridge this gap by providing both empirical findings and S&P implications for critical stakeholders at play.
@@ 3 Methods
Recruitment. In July 2023, we advertised our study on VR-related Discord servers, sub-Reddits, LinkedIn, and Facebook groups. Prior to the advertisement, we obtained permission from all the channels’ moderators, admins, or owners. In addition, one of the researchers reached out to a local school via personal connections. After we obtained the approval from the school, we advertised through the school’s official email communications. When advertising our study, we did not specify our S&P focus because we did not want to prompt our participants prior to the interview. We were interested in what concerns would first occur to parents in a natural response without preparation beforehand. Similar to Kumar et al.’s [16] approach, we framed it as an investigation into parents’ perceptions and practices of children’s VR usage. In our advertisement posts and emails, we attached a link to the pre-screening survey and asked all interested candidates to fill out the survey.
@@
From July to September 2023, we conducted one-hour remote interviews via Zoom. To address our RQs, each interview consisted of four major sections (see the full list of interview questions in Appendix A).
@@
First, to have a better context building on their pre-screening responses, we asked parents to have a more detailed account of their children’s and their VR usage. Then, we asked them about their perceptions of the general risks and benefits of children’s VR usage. We did not ask participants about their S&P-specific perceptions and practices until the next section, because we wanted to see if they would raise any S&P risks on their own initiative. If not, we were curious about their primary concerns.
@@
Following this section, we explicitly informed parents of our S&P focus and asked them to specify any S&P concerns they had. Additionally, we asked parents about their perceived stakeholders and each stakeholder’s responsibilities in addressing their concerns.
@@
Last, we asked parents what strategies they had adopted to address the concerns, if at all. We also asked them about their satisfaction with the existing S&P controls in VR and their desired features in future VR platforms. We closed the interviews with a live survey about parents’ awareness and perceived (or anticipated) effectiveness of the existing parental controls in VR. The live survey questions are in Appendix A. Before parents filled out the survey, we encouraged them to express their thoughts as they looked at the questions. For all the participants who completed the interviews, we provided one 20 USD Amazon gift card.
@@
Pilot interviews. Prior to the formal interviews, we conducted two pilot interviews with external researchers with expertise in social VR moderation, privacy, and security. We made sure our interview could be completed within one hour and asked them for advice on our interview protocol. In the pilot interviews, our live survey was a post-interview survey that two participants completed after the interviews ended. However, we found that a live survey using the quiz function in Zoom could be more useful in prompting participants’ holistic and nuanced thoughts on the features. Therefore, instead of using a post-interview survey, we switched to a live survey as a think-aloud protocol [74].
@@
Data collection and analysis. We hosted the screening and compensation surveys using Qualtrics [75]. All the interviews were audio-recorded and transcribed in English using the auto-transcription service provided by Zoom. One researcher carefully checked the accuracy of each transcript prior to the start of the analysis.
@@
Following Saldaña’s [76] coding approach, we conducted an in-depth thematic analysis of our interview transcripts. First, two researchers coded three transcripts independently. After discussing and incorporating the codes, we created an initial codebook that the two researchers agreed on. Using the initial codebook, one of the researchers coded ten of the remaining transcripts, while the other coded the other seven. The two researchers discussed their codes and resolved any conflicts through several weekly meetings. Since the codebook was generated after multiple discussions, inter-rater reliability was not necessary to be checked [77].
@@
We created codes about parents’ concerns for children’s VR activity (e.g., privacy-related, physical health), parents’ mitigation strategies (e.g., proactive verbal education), and parents’ perceived responsible stakeholders (e.g., VR companies, parents). Then, we merged similar codes and categorized them into emerging themes and sub-themes. Eventually, our codebook consisted of 13 themes, 95 sub-themes, and 613 codes.
@@
As is typical with qualitative studies, our recruited sample size was relatively small. The study relied on participants’ self-reported information, which may be subject to social desirability bias. As a consequence, our findings may not be generalizable to all parents of children who use VR. To mitigate this issue, our recruitment was on a rolling basis as mentioned earlier. We summarized the key insights after each interview and believed that after the sixteenth interview, participants’ insights reached the thematic saturation [78], i.e., no more new concepts or themes emerged. We then stopped inviting more candidates other than the already scheduled ones. In doing so, we attempted to cover as many diversified perspectives as possible for an interview study.
@@ 4 Findings
We begin this section by reporting the demographics and VR usage of our participants. Then, we report on parents’ perceptions of children’s S&P in VR (RQ1). We then elucidate parents’ different mitigation strategies for children’s S&P in VR (RQ2). Lastly, we show parents’ desired S&P-enhancing features and actions in VR from different stakeholders (RQ3).
@@
Although we did not require participants to specify their occupations, many of them disclosed the industries where they worked during the interviews, especially those working with VR/AR technologies and valuing S&P of technologies in use. Among the 15 participants who disclosed their professions, seven participants worked in technology companies, many of which focused on VR/AR technology and/or valued the S&P of their products. Two participants worked in finance and also had some expertise in S&P. Three participants were K12 educators who had used VR as teaching materials or worked in education technology companies. One participant was a game streamer with a lot of experience with VR in general and social VR in particular. The varied experiences of our participants provided us with rich and diversified insights into how parents from different industries and backgrounds approached their children’s S&P in VR. The VR usage details of our study participants and their children are included in Tables II and III in Appendix B. The full demographic details are specified in Table IV in Appendix C.
@@
In-home VR ownership and parents’ usage. We only interviewed parents who had VR devices in the household. We asked participants to self-describe the VR devices they own in the screening survey and verified the ownership at the beginning of the interview (see Table II in Appendix B). Almost all participants (17) only owned one VR headset shared with other family members, including children.
In the screening survey, parents chose their purposes of usage as gaming (16), creativity (8), socialization (8), education (8), streaming and films (7), health (6), and productivity/work (4). During the interviews, we also asked participants how frequently they used VR. Seven participants reported their frequency to be on a daily basis, eight participants were on a weekly basis, three participants were on a quarter basis, and two participants were not actively using VR at the time of the interview but indicated their children were more active users.
@@
As shown in Table III in Appendix B, according to the parents, their children’s purposes of VR usage included gaming (15), creativity (3), education (4), socialization (2), health (2), streaming and watching movies (1), others or unspecified (5).
Parents’ and children’s socialization in VR Particularly, since many S&P risks in VR currently take place during social interaction [13, 36, 37], we also asked participants to specify their socialization experiences in VR. 16 participants had at least tried social VR applications or connected with others in VR. Ten participants only met people they already knew in real life, including friends and family members. Six participants socialized with strangers in VR. Notably, one participant was a VR streamer who regularly met many new people in social VR.
@@
These details about parents’ and children’s VR usage are crucial to understanding parents’ awareness, attitudes, and coping strategies toward children’s S&P risks in VR, which will be detailed in the remainder of this paper.

### v3

@@ Abstract
Recent years have seen a sharp increase in the number of underage users in virtual reality (VR), where security and privacy (S&P) risks such as data surveillance and self-disclosure in social interaction have been increasingly prominent. Prior work shows children largely rely on parents to mitigate S&P risks in their technology use. Therefore, understanding parents’ S&P knowledge, perceptions, and practices is critical for identifying the gaps for parents, technology designers, and policymakers to enhance children’s S&P. While such empirical knowledge is substantial in other consumer technologies, it remains largely unknown in the context of VR. To address the gap, we conducted in-depth semi-structured interviews with 20 parents of children under the age of 18 who use VR at home. Our findings highlight parents generally lack S&P awareness due to the perception that VR is still in its infancy. To protect their children’s interactions with VR, parents currently primarily rely on active strategies such as verbal education about S&P. Passive strategies such as using parental controls in VR are not commonly used among our interviewees, mainly due to their perceived technical constraints. Parents also highlight that a multi-stakeholder ecosystem must be established towards more S&P support for children in VR. Based on the findings, we propose actionable S&P recommendations for critical stakeholders, including parents, educators, VR companies, and governments.
@@ 1 Introduction
Furthermore, users including children have flooded social VR applications (e.g., Horizon Worlds, Rec Room, and VRChat) [5, 3, 4], raising S&P concerns. In internet-enabled social VR applications, besides data surveillance, users also face S&P risks in social interaction, including self-disclosure [13]. Prior work has suggested that social VR users easily disclose a wide range of personal information to strangers, including emotions, life experiences, voices, and even real-life appearances [13].
@@ 2 Related Work
S&P risks in VR. VR technologies collect a vast amount of sensitive data that has been categorized by Dick [7] into four types – (1) observable data that generate users’ virtual presence, e.g., avatars and real-time in-world interactions [28, 29]; (2) observed data that enhance immersive experiences, e.g., motion/eye-tracking [8, 9, 29, 30, 31]; (3) computed data that improve user services, e.g., user profiles for advertisements; and (4) associated data that allow Internet-enabled functions, e.g., IP addresses, usernames, and friend lists [32]. Unlike many other consumer technologies, VR collects these sensitive data that are indispensable for its core utility – immersive experiences [33]. Specifically, users have to trade their motion-tracking data for the full-body tracking experience; the data collected via external-facing cameras for detecting obstacles around; and voice data for social experiences [7]; just to name a few. Due to the unique immersive nature of VR technologies, researchers have been calling for policymakers to reform existing privacy laws and policies such as COPPA [34] and HIPAA [35] that would unnecessarily limit users’ experiences of VR technologies [7]. In addition, a 2018 study has revealed that VR developers considered VR privacy policies problematic for providing developers with limited standards and guidelines regarding S&P [6].
@@
In the context of social interaction, more S&P risks have emerged as VR evolves to become a popular form of social technology [13, 36, 37]. Mainstream social VR platforms, including VRChat, Horizon Worlds, Rec Room, and more, allow users to socialize with others through head-mounted displays (HMDs) [38, 39]. In this vein, a volume of social VR work has identified numerous privacy and safety risks. Maloney et al. [13] have found that social VR users tend to feel comfortable and safe to share emotions and life experiences due to the anonymity afforded by social VR. In addition, users’ physical traits, such as their voices and real-life appearances, can also be disclosed since (1) voice chats are ubiquitous in social VR and (2) some users prefer to represent themselves with a virtual avatar that looks similar to their appearance in real life. More recently, with the increasing popularity of social VR streaming, users’ privacy has been violated by streamers who broadcast in public VR spaces [36].
In systematizing VR threats from the existing literature, a recent SoK paper [40] has characterized adversary types corresponding to four data-processing entities in VR – hardware, client, server, and user adversaries. Hardware adversaries can access users’ raw sensor data and manipulate information provided to VR applications and users. Client adversaries characterize how developers of VR applications and content creators may convey “misinformative, manipulative, and deceptive content” to users. Server adversaries can enable multi-user functionality and process networked data presented to users. Finally, user adversaries represent other users in the same VR application. After characterizing this threat model, Garrido et al. [40] have pointed out that user adversarial attacks easily occur, whereas hardware adversaries are still not mature. Importantly, due to the lack of privacy policy standards and enforcement [41], server adversaries are also easily executable, inferring many sensitive data such as user demographics in a short time.
@@
VR risks for children. Due to the age-mixed environment [4, 3, 42], children who are considered part of the marginalized community in VR often face more tailored risks, such as predators, harassment, cyberbullying, inappropriate adult content [5, 43, 44, 45, 46], etc. However, compared to such safety threats that have been widely studied, children’s S&P in social VR have been much more marginalized in scholarly work. Previous studies have collectively pointed out that bystanders and parents have concerns for children’s privacy as they interact with strangers in VR [4, 3]. Notably, one study has suggested VR privacy guidelines be made more straightforward to align with children’s reading and literacy levels [5]. To further provide actionable S&P-enhancing recommendations, more research centered on children’s S&P is needed. Since children heavily rely on parents’ decision-making of their online S&P [14, 47, 15, 48, 16], a critical gap exists in terms of parents’ attitudes and knowledge around children’s S&P in VR.
@@
Parents’ risk mitigation strategies toward technologies at home. Nissenbaum defines privacy as “neither a right to secrecy nor a right to control, but a right to appropriate flow of personal information.” [49]. This definition is especially applicable to online social contexts where privacy risks arise from self-disclosure [50, 49, 51, 52, 53, 54]. When self-disclosure occurs online, privacy is about the appropriate flow of information [55]. With children’s significantly increased online activities [56, 57, 58], more unexpected information flows have raised new privacy concerns [16, 59]. For instance, information disclosed via email can be retained longer than face-to-face communication [16]. It is critical to protect children’s privacy online since (1) privacy is an important space for children to practice their decision-making, create boundaries, and foster independence [50, 55]; (2) privacy can benefit children’s experience in identity play, relationship building, and peer communication [60, 16].
@@
Security is another important aspect of children’s interactions with technology. Prior work suggests three goals of information security [16, 61]: (1) maintaining the integrity of personal information stored and transmitted through networks, (2) authenticating users with access to personal information, and (3) preserving the confidentiality of personal information. This work uses the above-mentioned definitions of S&P to unpack parents’ risk perceptions and mitigation strategies.
@@
Given the importance of S&P for children’s technology use, parents have the responsibility in protecting children’s digital S&P at home [62, 63]. Kumar et al. [16] have found that children aged 9 years old or younger are most likely to seek help from parents when confronted with unknown risks online and largely rely on parents’ explicit rules to deal with their digital S&P risks. Other research in this vein has suggested that children of different ages trust their parents’ S&P protections and guidance such as on data sharing [64, 65, 62]. Additionally, another line of research has suggested that parents’ S&P attitudes and practices may shape that of children’s [62, 66, 21, 67, 17, 18, 19]. For example, Shin et al. [17] have reported that parents who self-identify as more S&P-focused also tend to better direct children’s S&P behavior online.
@@
Parents adopt a variety of strategies to mitigate the risks of children’s technology use [16], including talking to children, viewing children’s accounts, using parental controls, being around their children during their technology use, setting rules and boundaries, etc. [68, 69, 20, 21]. Having examined different parental strategies, Wisniewski et al. [22] have summarized three types of strategies in the Teen Online Safety Strategies (TOSS) framework: (1) passive monitoring, (2) rules and restrictions in usage, and (3) active mediation such as through having conversations with children. However, when mitigating the risks, parents face a number of challenges, such as underestimating the potential risks faced by children [70, 71, 23], undermining parent-child trust [23, 72, 73, 74], and an incomplete understanding of technologies compared to their children [23, 75, 24]. Therefore, it is crucial to investigate parents’ S&P practices around children’s technology use at home and identify the tradeoffs in their decision-making to help parents and technology designers work together towards safer and more inclusive technologies for children.
@@ 3 Methods
Recruitment. In July 2023, we advertised our study on VR-related Discord servers, sub-Reddits, LinkedIn, and Facebook groups. Prior to the advertisement, we obtained permission from all the channels’ moderators, admins, or owners. In addition, one of the researchers reached out to a local school via personal connections. After we obtained the approval from the school, we advertised through the school’s official email communication. When advertising our study, we did not specify our S&P focus because we did not want to prompt our participants prior to the interview. We were interested in what concerns would first occur to parents in a natural response without preparation beforehand. Similar to Kumar et al.’s [16] approach, we framed it as an investigation into parents’ perceptions and practices of children’s VR usage. In our advertisement posts and emails, we attached a link to the pre-screening survey and asked all interested candidates to fill out the survey.
@@
From July to September 2023, we conducted one-hour remote interviews via Zoom. To address our RQs, each interview consisted of four major sections.
@@
First, to have a better context building on their pre-screening responses, we asked participants to have a more detailed account of their children’s and their VR usage. Then, we asked them about their perceptions of the general risks and benefits of children’s VR usage. We did not ask participants about their S&P-specific perceptions and practices until the next section, because we wanted to see if they would raise any S&P risks on their own initiative. If not, we were curious about their primary concerns.
@@
Following this section, we explicitly informed participants of our S&P focus and asked them to specify any S&P concerns they had. Additionally, we asked participants about their perceived stakeholders and each stakeholder’s responsibilities in addressing their concerns.
@@
Last, we asked participants what strategies they had adopted to address the concerns, if at all. We also asked them about their satisfaction with the existing S&P controls in VR and their desired features in future VR platforms. We closed the interviews with a live survey about participants’ awareness and perceived (or anticipated) effectiveness of the existing parental controls in VR. Before participants filled out the survey, we encouraged them to express their thoughts as they looked at the questions. For all the participants who completed the interviews, we provided one 20 USD Amazon gift card.
@@
Pilot interviews. Prior to the formal interviews, we conducted two pilot interviews to assess the interview duration and clarity of interview questions [76]. The two pilot participants were external researchers with expertise in VR moderation, privacy, and security. One of them was a parent. The pilot participants’ relevant research expertise and parental perspectives helped us fine-tune our interview questions. We made sure our interview could be completed within one hour and asked them for advice on our interview protocol. When using the revised protocol in formal interviews, no participants expressed any confusion, indicating the success of pilot interviews. After the pilot interviews, we employed a post-interview survey. However, we found that a live survey using the quiz function in Zoom could be more interactive and useful in eliciting participants’ holistic and nuanced thoughts on the features. Therefore, instead of using a post-interview survey, we switched to a live survey as a think-aloud protocol [77].
@@
Data collection and analysis. We hosted the pre-screening and compensation surveys using Qualtrics [78]. All the interviews were audio-recorded and transcribed in English using the auto-transcription service provided by Zoom. One researcher carefully checked the accuracy of each transcript prior to the start of the analysis.
@@
Following Saldaña’s [79] coding approach, we conducted an in-depth thematic analysis of our interview transcripts. First, two researchers coded three transcripts independently. After discussing and incorporating the codes, we created an initial codebook that the two researchers agreed on. Using the initial codebook, one of the researchers coded ten of the remaining transcripts, while the other coded the other seven. Similar to many other qualitative S&P studies [80, 81, 82], the two researchers discussed the codes and resolved any conflicts through several weekly meetings. During our analysis, as we prioritized our primary goal to generate emerging themes and concepts, we did not calculate the inter-rater reliability (IRR) to seek theoretical agreement [83].
@@
We created codes about parents’ concerns for children’s VR activity (e.g., privacy-related, physical health), parents’ mitigation strategies (e.g., proactive verbal education), and parents’ perceived responsible stakeholders (e.g., VR companies, parents). Then, we merged similar codes and categorized them into emerging themes and sub-themes. Eventually, our codebook consisted of 13 themes, 95 sub-themes, and 613 codes. All the study artifacts, including the final codebook, can be retrieved from Appendix B.
@@
As is typical with qualitative studies, our recruited sample size was relatively small. The study relied on participants’ self-reported information, which may be subject to social desirability bias. In addition, we did not calculate our IRR. As a consequence, our findings may not be generalizable to all parents of children who use VR. To mitigate this issue, our recruitment was on a rolling basis as mentioned earlier. We summarized the key insights after each interview. Based on the guidelines from prior work [84, 85], we identified data saturation after the sixteenth interview, i.e., no more new insights emerged. We then stopped inviting more candidates other than the ones that had already been scheduled. As a result, we conducted four more interviews after identifying saturation, in line with recommended practices [85]. In doing so, we attempted to cover as many diversified perspectives as possible for an interview study.
@@ 4 Findings
Although we did not require participants to specify their occupations, many of them disclosed the industries where they worked at the time of the interview, especially those working with VR/AR technologies and valuing S&P of technologies. Among the 15 participants who disclosed their professions, seven participants worked in technology companies, many of which focused on VR/AR technology and/or valued the S&P of their products. Two participants worked in finance and also had some expertise in S&P. Three participants were K12 educators who had used VR as teaching materials or worked in education technology companies. One participant was a game streamer with abundant experience in VR. The varied experiences of our participants provided us with rich and diversified insights into how parents from different industries and backgrounds approached their children’s S&P in VR. The VR usage details of our study participants and their children are included in Table II in Appendix A. The full demographic details are specified in Table III in Appendix A.
@@
In-home VR ownership and parents’ usage. We only interviewed parents who had VR devices in the household. We asked participants to self-describe the VR devices they own in the pre-screening survey and verified the ownership at the beginning of the interview (see Table II in Appendix A). Almost all participants (17) only owned one VR headset shared with other family members, including children.
In the pre-screening survey, parents selected their purposes of usage as gaming (16), creativity (8), socialization (8), education (8), streaming and films (7), health (6), and productivity/work (4). During the interviews, we also asked participants how frequently they used VR. Seven participants reported their frequency to be on a daily basis, eight participants were on a weekly basis, three participants were on a quarter basis, and two participants were not actively using VR at the time of the interview but indicated their children were more active users.
@@
As children of different ages may face different risks in VR, we paid attention to how the age of children may impact parents’ perceptions and practices that are closely tied to children’s S&P in VR. Therefore, in our findings, we frequently compare parents of children above and below the recommended age for usage, focusing on their different S&P concerns and strategies, if any.
@@
As shown in Table II in Appendix A, according to the parents, their children’s purposes of VR usage included gaming (15), creativity (3), education (4), socialization (2), health (2), streaming and watching movies (1), others or unspecified (5).
Parents’ and children’s socialization in VR Particularly, since many S&P risks in VR currently take place during social interaction [13, 36, 37], we also asked participants to specify their socialization experience in VR. 16 participants had at least tried social VR applications or connected with others in VR. Ten participants only met people they already knew in real life, including friends and family members. Six participants socialized with strangers in VR. Notably, one participant was a VR streamer who regularly met many new people in social VR.
@@
These details about parents’ and children’s VR usage are crucial in understanding parents’ awareness, attitudes, and coping strategies toward children’s S&P risks in VR, which will be detailed in the remainder of this paper.